Calculator the acceleration on a 2.5 x 10^6 kg rocket if the net thrust of the rocket 2.8 x 10^7 N.

1
Expert's answer
2022-04-12T16:12:46-0400
"a=\frac{F}{m}=\frac{2.8*10^7}{2.5*10^6}=11.2\;\rm m/s^2"